OpenStudy (angela210793):

10 years ago A bought a share for 20$ and sold it today for 35$.What is the rate of profit he had from this investment. I thought to use PV=FV*FVIF where PV=35 and PV=20 and then calculate FVIF..is this right?

OpenStudy (angela210793):

hang on a sec

OpenStudy (angela210793):

I don't get it O.o :(

OpenStudy (angela210793):

Annam terms? what's that?

OpenStudy (angela210793):

so its PVA?

OpenStudy (anonymous):

no

OpenStudy (angela210793):

O.o What is it then? O.o

OpenStudy (angela210793):

the return today i guess

OpenStudy (angela210793):

ok...then what?

OpenStudy (angela210793):

:O how's that??? we've never solved a problem in this way O.o :( I'm gonna fail :(((((

OpenStudy (anonymous):

oh in this question is discounting rate given ???

OpenStudy (angela210793):

no...i wrote it exactly as it is written in my w-book....

OpenStudy (angela210793):

is there any way solving this using PV or FV or anything like this?

OpenStudy (angela210793):

ok ok

OpenStudy (angela210793):

:( ok thanks anyway :)))

OpenStudy (anonymous):

see the thing which u were saying was including the interest factor right Future value = Present Value *(1 + K)^n 35 = 20*(1+k)^10 on solving this one u will get k = 0.0575 Rate of return = 5.75 %

OpenStudy (angela210793):

I actually had this solved in my notebook and my teacher had solved this using PV=FV*PVIF

OpenStudy (anonymous):

Future value = Present Value *(1 + K)^n 35 = 20*(1+k)^10 on solving this one u will get k = 0.0575 Rate of return = 5.75 %

OpenStudy (anonymous):

first i had taken the layman's approach and in that i was ignoring the discounting factor

OpenStudy (anonymous):

hey hang on you have written PV = FV * PVIF is that so

OpenStudy (angela210793):

yea O.o

OpenStudy (anonymous):

ok PV = ?? 20 or, 35

OpenStudy (angela210793):

PV=20 FV=35 PVIF is wht i have to find...which means i waas right except the fact tht instead of writing PVIF i wrote FVIF

OpenStudy (angela210793):

yea yea :) thanks :) going to eat now :) cya :)

OpenStudy (anonymous):

^^^^^ see above i had solved that using discounting factor too above